The reaction of diatomic fluorine gas (F2) with compound X yields a single product, compound Y. The mass percent composition of the product, compound Y is 61.7% F2 and 38.3% Cl. What is the probable empirical formula for product Y? What is a probable empirical formula for compound X?

1
Expert's answer
2022-05-19T03:13:03-0400

"F_aCl_b"

"w(F)=\frac{aM(F)x100}{aM(F)+bM(Cl)}=61.7"

"w(Cl)=\frac{bM(Cl)x100}{aM(F)+bM(Cl)}=38.3"

M(F)=19 g/mole

M(Cl)=35.5 g/mole

"\frac{a(19)100}{19a+35.5b}=61.7"

"1900a=1172.3a+2190.35b"

"b=\frac{a(1900-1172.3)}{2190.35}=0.33a"

a=3b

empirical formula for product Y is"F_3Cl"

"3F_2+X \rightarrow 2F_3Cl"

Then X is Cl2